# Transient QPOs of Fermi-LAT blazars with Linearly Multiplicative Oscillations

### v1

@@ 4 Theoretical Models
Shock waves can also be generated by the interaction of external agents with the jet. One example is the penetration of dense clouds from the broad-line region into the jet. As these clouds interact with the relativistic flow, they induce strong shocks that efficiently accelerate particles, leading to enhanced synchrotron and inverse Compton emission, particularly in the -ray band (del Palacio et al., 2019). Another source of shocks can be the collision between the jet and stellar winds from massive stars within the host galaxy. These winds inject additional turbulence and inhomogeneity into the jet, generating shocks that modify the particle acceleration processes and impact the emission (Bicknell & Wagner, 2002).
@@
Another phenomenon that can offer a compelling explanation for the dynamic evolution of the oscillation amplitude is the precession of the jet. In the scenario, changes in the Doppler boosting due to the variable orientation of the jet with respect to the line of sight cause changes in the amplitude of the oscillations over time (Camenzind & Krockenberger, 1992). This is particularly evident when the precessing jet crosses critical angles that either amplify or attenuate the emitted radiation, creating a distinct trend in the LC (e.g., Chen & Jiang, 2024). The precession could be originated by multiple scenarios. One scenario could be the Lense-Thirring precession (e.g., Franchini et al., 2016; Zanazzi & Lai, 2019). This phenomenon is a relativistic effect caused by the dragging of spacetime due to the rotation of an SMBH. In blazars, the jet is typically aligned with the accretion disk’s angular momentum axis. However, if the accretion disk is misaligned with the SMBH spin axis, the frame-dragging effect from the rotating SMBH induces precession in the inner regions of the disk. This Lense-Thirring precession gradually changes the orientation of the jet, leading to observable variations in the direction of the jet over time. Alternatively, the presence of a binary SMBH system at the center of the blazar can be the origin of such precession (e.g., Graham et al., 2015; Qian, 2022). In the binary scenario, gravitational torques within the SMBH system induce precession, causing the relativistic jet of one SMBH to follow a slow, precessional path influenced by its companion. This gradual angular shift changes the jet angle relative to our line of sight, modifying the Doppler boosting effect (e.g., Villata & Raiteri, 1999).
@@ 4.1 Interpretation
In this paper, we interpret the observed QPOs with multiplicative amplitudes as the result of a combination of different phenomena. First, we consider that QPOs arise from a plasmoid following a helical trajectory dictated by the magnetic field geometry. As discussed in 4, this motion alters the angle of our line of sight with respect to the jet, leading to periodic variations in the Doppler boosting factor. Additionally, the movement of the plasmoid can generate shock waves, which, in turn, induce fluctuations in the emitted flux. This effect can be perceived as an apparent change in the jet’s axis and, consequently, in our viewing angle (Wiita, 2011), modulating the overall flux observed from the jet.
Combined with this, we also use the conclusions presented in Rieger (2004) to explore how differential Doppler boosting, caused by the helical motion of emitting components within the jet, can lead to observable periodicity. This motion is nonballistic, meaning the radiating plasma elements follow curved trajectories rather than moving in straight lines. Rieger (2004) identified three possible physical mechanisms for helical jet motion in blazars. The first is nonballistic motion driven by the orbital motion in an SMBH binary. The second involves either ballistic or nonballistic helical jet paths induced by jet precession, which may result from gravitational torques exerted by a secondary black hole or from accretion disk warping. The third mechanism attributes nonballistic helical motion to intrinsic jet instabilities. The first two scenarios are particularly viable for explaining 1 year.
@@ 4.1.1 Model expressions
as presented in Otero-Santos et al. (2023). Additionally, the emitted flux can be expressed as
@@ 4.1.2 Flux Fit
Examining the shape of the oscillations (see Figure 1), we observe that their amplitude initially increases over time, reaching a peak, and subsequently decreases. Consequently, we substitute in Equation (4) with a linear function as =. This approach suggests that the amplitudes evolve over time through a multiplicative relationship (Peñil et al., 2025c). Therefore, our model accounts for the modulation introduced by the impact of the shock wave on the plasmoid’s emission. To extract the linear component of the flux, we employ a linear regression222We use the LinearRegression function of the Python package Scikit-learn, which is optimized specifically for linear fitting., characterizing the parameters of the linear trend333We use the seasonal_decompose function of the Python package Statsmodels..
@@
We then test the hypothesis of the LC being accurately described by this model through the R2 parameter, a statistical measure that assesses the goodness of fit of a regression model. The value of R2 ranges from 0 to 1 (or 0 to 100 in percentage), where the higher values denote a better fit and suggest that the model can better explain the variations in the data. In Hair et al. (2011), three different levels of R2—25%, 50%, and 75%—are outlined and categorized as weak, moderate, and substantial, respectively to evaluate the accuracy of a fit.
@@ 5 Results and Discussion
The Lorentz factor, , generally falls within the range of approximately 10 to 20 for FSRQs (Ghisellini et al., 2010), often reaching values (Xiong & Zhang, 2014). In this study, we adopt (Ghisellini & Tavecchio, 2009; Sarkar et al., 2021). The parameter is directly related to through the expression , which yields . The pitch angle, , is another variable with a potentially broad range, typically between and (e.g., Jorstad et al., 2017; Butuzova, 2018). However, owing to the lack of observational constraints in the literature, we adopt a pitch angle value of as done in previous studies (Sarkar et al., 2021; Prince et al., 2023).
For the values of the period , the linear slope, the intrinsic spectral index , and the Doppler boosting factor , we conduct an exploratory process to achieve the optimal flux fit modeled by the Equation (4) for each LC by maximizing the R2 value. For the periods and slopes, this optimization spans the range of periods and slopes determined through the methodology outlined in . The resulting values are summarized in Table 2.
Regarding , Singal et al. (2014) conducted a comprehensive analysis of blazars detected by Fermi-LAT, reporting a mean photon spectral index of 2.4. Consistently, we explore values for in the range of 1.2, 1.3, 1.4, 1.5, and 1.6 to better understand the variations in the spectral behavior.
Finally, for the Doppler boosting factor , we explore different values, from to . When the emission in the jet is primarily due to synchrotron radiation, . This lower value reflects the isotropic nature of synchrotron emission in the rest frame of the jet, typically prominent in radio to optical wavelengths and causing moderate flux variability in these bands (Urry & Padovani, 1995). For rays, which are often produced via inverse Compton scattering, tends to be higher, typically 3 or 4 (Dermer, 1995). This increase occurs because inverse Compton scattering — and especially external Compton scattering, where seed photons from outside the jet are scattered — is more sensitive to the Doppler factor. This phenomenon is particularly dominant in FSRQs, which constitute the objects in our sample. Finally, is particularly relevant in scenarios where short-lived or transient events are observed, suggesting that a localized, highly energetic region within the jet contributes to extreme, rapid flux variations characteristic of the FSRQ (Dermer, 1995). The specific values of the periods, slopes, , and that maximize the R2 value are displayed in Figure 2 and Table 2, highlighting the most accurate fits achieved for each case.
@@
As discussed in 4.1.2, we perform a detailed analysis of the QPOs by separately examining the increasing and decreasing segments of the LC, characterizing potential variations in amplitude or period between the rising and falling phases. We apply Equation (4) to model each segment, optimizing the variable values to maximize R2.
@@
To illustrate the evolution of , we present Figure 3, which depicts its temporal variation based on different parametrization results for the three blazars. This visualization shows how evolves over time, helping to assess the theoretical scenarios considered for interpreting the observed flux variations. In the specific case of 4C +31.03, we observe that the viewing angle evolution shows a discontinuity of 30% between the increasing and decreasing segments of the LC. Both the low R2 value and this discontinuity suggest potential inconsistencies in modeling the decreasing segment. The heightened variability within this portion of the LC could be impacting the analysis accuracy, as fluctuations in the -ray emission introduce complexities that may not align well with the current model assumptions. These results raise questions about the suitability of the selected theoretical model in capturing the transient QPO behavior of 4C +31.03, suggesting that an alternative approach may be necessary to better understand the emission mechanisms in this blazar. One possible scenario is that recurring boosts in the turbulent flow behind a propagating shock generate the QPO. Due to the stochastic nature of turbulence, variations in its properties can cause fluctuations in both the QPO period and the emitted flux (Wiita, 2011), leading to changes in the observed period, as seen in the case of 4C +31.03 (see Table 2).
@@ 5.2 MG1 J123931+0443
The flux fits for both segments of the LC separated by the flare located at 56400 MJD present consistent results, with compatible periods and trends observed across both segments, as shown in Table 2 and Figure 2. Additionally, Figure 3 reveals no significant discontinuity, with only a minor variation of 6.6%, which can be attributed to uncertainties introduced by the flaring state during parameter estimation. In fact, from the QPO analysis, we observe this error in the is consistent within errors at a 1 confidence level. These findings reinforce the accuracy of the chosen emission model, supporting it as a plausible scenario to explain the QPO observed in MG1 J123931+0443.
As discussed in 4, several theoretical scenarios can provide explanations for the observed QPO. Rieger (2004) suggests that periodicity driven by differential Doppler boosting could result from (1) orbital motion in a close binary SMBH, (2) Newtonian precession within a close SMBH, or (3) internal jet rotation. Among these, Newtonian-driven precession appears particularly viable for explaining periods of year, which aligns with the observed period in MG1 J123931+0443, 285 days. If the period observed in this object is associated with the presence of a binary SMBH, we can relate with the intrinsic orbital period (), as
where is the redshift of the source, and the motion of the plasma is considered as nonballistic and helical, driven by the binary system’s orbital dynamics (Rieger, 2004). A nonballistic motion occurs when the trajectory of the plasma is continuously influenced by external forces, such as the jet’s environment or the central engine of the blazar (Rieger, 2004). For an observed period days and (see Table 1), the calculated orbital period is years assuming a .
Using this orbital period, we estimate the mass of the primary SMBH, , based on the formulation provided by Ostorero et al. (2004),
where the mass ratio typically ranges from to (Ostorero et al., 2004). Substituting the orbital period, we estimate the primary SMBH’s mass to range from 7.7 to 1.2, consistent with reasonable mass estimates for SMBHs (Liang & Liu, 2003). Pacciani et al. (2012) estimated the SMBH mass of this blazar using various techniques, obtaining a range of 2.1-9.7, consistent with the range of values reported here.
@@
is the gravitational constant and is the mass of the secondary SMBHs. This yields an orbital separation in the range of 8.9 to 1.8 parsecs.
If the primary SMBH emits a relativistic jet, its behavior may be influenced by the binary system’s orbital dynamics. The orbital radius, defined as the distance between the primary SMBH and the system’s barycenter (Ostorero et al., 2004), is calculated as:
@@ 5.3 PKS 1622253
This blazar exhibits a high level of consistency between the estimated period and trend for both segments of the LC, further supported by the minimal discontinuity in , measured at 5%. The observed period, days, also suggests that Newtonian precession due to the interaction within a binary SMBH system could be a plausible physical scenario, as in the case of MG1 J123931+0443.
We characterize the candidate binary SMBH system of PKS 1622253. Using the redshift value of and the aforementioned expressions. We derive an orbital period of 117 years. The estimation of the primary SMBH’s mass ranges from 2.0 to 2.2, consistent with the mass of 4.0 reported by Ghisellini et al. (2010). The orbital separation ranges of 1.8 to 3.7 parsecs. The orbital radius ranges between 9.2 and 3.7 parsecs.
@@ 5.4 Precession Model
An alternative scenario could be the precessing jet model, which was mentioned in 4. In this scenario, the relativistic jet emitted by a blazar undergoes precession, causing the jet’s orientation to change over time. This variation alters the angle between the jet and our line of sight, influencing the observed emission due to changes in the relativistic beaming effects, potentially leading to the observed periodicity in the blazar’s LCs (Camenzind & Krockenberger, 1992).
@@ 6 Summary
Our findings indicate that the QPO characteristics observed in PKS 1622253 and MG1 J123931+0443 could be consistent with the predictions of a binary SMBH system, suggesting that the periodic variability might be explained by precessional dynamics associated with a binary SMBH system. However, given the relatively low significance derived from our transient QPO analysis, these results should be interpreted cautiously as tentative hints rather than definitive evidence. Additional studies and observations will be essential to further test and validate this interpretation. Despite these limitations, the consistency between the theoretical model and the observed -ray LCs highlights the plausibility that binary SMBH systems could be involved in producing QPOs in blazars, offering potentially valuable insights into the nature of periodic variability in these sources.
@@ Acknowledgements.
P.P. and M.A. acknowledge funding under NASA contract 80NSSC20K1562. J.O.-S. acknowledges founding from the Istituto Nazionale di Fisica Nucleare (INFN) Cap. U.1.01.01.01.009. This work was supported by the European Research Council, ERC Starting grant MessMapp, S.B. Principal Investigator, under contract no. 949555, and by the German Science Foundation DFG, research grant “Relativistic Jets in Active Galaxies” (FOR 5195, grant No. 443220636).References
- Abdo et al. (2010) Abdo A. A., et al., 2010, ApJ, 722, 520
- Adhikari et al. (2024) Adhikari S., Peñil P., Westernacher-Schneider J. R., Domínguez A., Ajello M., Buson S., Rico A., Zrake J., 2024, ApJ, 965, 124
- Asada et al. (2008) Asada K., Inoue M., Nakamura M., Kameno S., Nagai H., 2008, The Astrophysical Journal, 682, 798
- Atwood et al. (2009) Atwood W. B., et al., 2009, ApJ, 697, 1071
- Ballet et al. (2020) Ballet J., Burnett T. H., Digel S. W., Lott B., 2020, arXiv e-prints, p. arXiv:2005.11208
- Bicknell & Wagner (2002) Bicknell G. V., Wagner S. J., 2002, PASA, 19, 129
- Butuzova (2018) Butuzova M. S., 2018, Astronomy Reports, 62, 116
- Camenzind & Krockenberger (1992) Camenzind M., Krockenberger M., 1992, A&A, 255, 59
- Chen & Jiang (2024) Chen Z.-H., Jiang Y., 2024, A&A, 689, A35
- Christie et al. (2019) Christie I. M., Petropoulou M., Sironi L., Giannios D., 2019, MNRAS, 482, 65
- Covino et al. (2018) Covino S., Sandrinelli A., Treves A., 2018, MNRAS, 482, 1270
- Czerny (2004) Czerny B., 2004, arXiv e-prints, pp astro–ph/0409254
- Dermer (1995) Dermer C. D., 1995, ApJ, 446, L63
- Dong et al. (2020) Dong L., Zhang H., Giannios D., 2020, MNRAS, 494, 1817
- Emmanoulopoulos et al. (2013) Emmanoulopoulos D., McHardy I. M., Papadakis I. E., 2013, MNRAS, 433, 907
- Fichet de Clairfontaine et al. (2022) Fichet de Clairfontaine G., Meliani Z., Zech A., 2022, A&A, 661, A54
- Finke (2024) Finke J. D., 2024, Frontiers in Astronomy and Space Sciences, 11, 1384234
- Franchini et al. (2016) Franchini A., Lodato G., Facchini S., 2016, MNRAS, 455, 1946
- Ghisellini & Tavecchio (2009) Ghisellini G., Tavecchio F., 2009, MNRAS, 397, 985
- Ghisellini et al. (2010) Ghisellini G., Tavecchio F., Foschini L., Ghirlanda G., Maraschi L., Celotti A., 2010, MNRAS, 402, 497
- Giannios (2013a) Giannios D., 2013a, MNRAS, 431, 355
- Giannios (2013b) Giannios D., 2013b, MNRAS, 431, 355
- Gracia et al. (2003) Gracia J., Peitz J., Keller C., Camenzind M., 2003, Monthly Notices of the Royal Astronomical Society, 344, 468
- Graham et al. (2015) Graham M. J., et al., 2015, Nature, 518, 74
- Greco et al. (2016) Greco G., et al., 2016, in Napolitano N. R., Longo G., Marconi M., Paolillo M., Iodice E., eds, Astrophysics and Space Science Proceedings Vol. 42, The Universe of Digital Sky Surveys. p. 105 (arXiv:1509.03342), doi:10.1007/978-3-319-19330-4˙16
- Hair et al. (2011) Hair J., Ringle C., Sarstedt M., 2011, Journal of Marketing Theory and Practice, 19, 139
- Heavens & Drury (1988) Heavens A. F., Drury L. O., 1988, Monthly Notices of the Royal Astronomical Society, 235, 997
- Jorstad et al. (2017) Jorstad S. G., et al., 2017, ApJ, 846, 98
- Liang & Liu (2003) Liang E. W., Liu H. T., 2003, MNRAS, 340, 632
- Lomb (1976) Lomb N. R., 1976, Ap&SS, 39, 447
- Marscher & Gear (1985) Marscher A. P., Gear W. K., 1985, ApJ, 298, 114
- Meyer et al. (2021) Meyer M., Petropoulou M., Christie I. M., 2021, ApJ, 912, 40
- Mohan & Mangalam (2015) Mohan P., Mangalam A., 2015, ApJ, 805, 91
- Nina Golyandina (2020) Nina Golyandina A. Z., 2020, Singular Spectrum Analysis for Time Series, 2 edn. S, Springer Berlin, Heidelberg, doi:https://doi.org/10.1007/978-3-662-62436-4
- Ostorero et al. (2004) Ostorero L., Villata M., Raiteri C. M., 2004, A&A, 419, 913
- Otero-Santos et al. (2023) Otero-Santos J., Peñil P., Acosta-Pulido J. A., Becerra González J., Raiteri C. M., Carnerero M. I., Villata M., 2023, MNRAS, 518, 5788
- Pacciani et al. (2012) Pacciani L., et al., 2012, MNRAS, 425, 2015
- Peñil et al. (2020) Peñil P., et al., 2020, ApJ, 896, 134
- Peñil et al. (2024a) Peñil P., et al., 2024a, MNRAS, 527, 10168
- Peñil et al. (2024b) Peñil P., et al., 2024b, MNRAS, 529, 1365
- Peñil et al. (2025a) Peñil P., Ajello M., Buson S., Domínguez A., Westernacher-Schneider J. R., Zrake J., 2025a, arXiv e-prints, p. arXiv:2211.01894
- Peñil et al. (2025b) Peñil P., Torres-Albà N., Rico A., Ajello M., Buson S., Adhikari S., 2025b, MNRAS, 539, 993
- Peñil et al. (2025c) Peñil P., Domínguez A., Buson S., Ajello M., Adhikari S., Rico A., 2025c, ApJ, 980, 38
- Penil et al. (2025) Penil P., Otero-Santos J., Banerjee A., Buson S., Rico A., Ajello M., Adhikari S., 2025, arXiv e-prints, p. arXiv:2507.03967
- Prince et al. (2023) Prince R., Banerjee A., Sharma A., Kumar das A., Gupta A. C., Bose D., 2023, A&A, 678, A100
- Qian (2022) Qian S. J., 2022, arXiv e-prints, p. arXiv:2206.14995
- Rico et al. (2024) Rico A., Domínguez A., Peñil P., Ajello M., Buson S., Adhikari S., Movahedifar M., 2024, arXiv e-prints, p. arXiv:2412.05812
- Rieger (2004) Rieger F. M., 2004, ApJ, 615, L5
- Rieger (2019) Rieger F. M., 2019, Galaxies, 7
- Ruan et al. (2012) Ruan J. J., et al., 2012, ApJ, 760, 51
- Sandrinelli et al. (2016) Sandrinelli A., Covino S., Dotti M., Treves A., 2016, AJ, 151, 54
- Sarkar et al. (2021) Sarkar A., Gupta A. C., Chitnis V. R., Wiita P. J., 2021, MNRAS, 501, 50
- Scargle (1982) Scargle J. D., 1982, ApJ, 263, 835
- Shukla & Mannheim (2020) Shukla A., Mannheim K., 2020, Nature Communications, 11
- Singal et al. (2014) Singal J., Ko A., Petrosian V., 2014, ApJ, 786, 109
- Sironi et al. (2016) Sironi L., Giannios D., Petropoulou M., 2016, MNRAS, 462, 48
- Sobacchi et al. (2017) Sobacchi E., Sormani M. C., Stamerra A., 2017, MNRAS, 465, 161
- Torrence & Compo (1998) Torrence C., Compo G. P., 1998, Bulletin of the American Meteorological Society, 79, 61
- Urry (2011) Urry M., 2011, Journal of Astrophysics and Astronomy, 32, 139
- Urry & Padovani (1995) Urry C. M., Padovani P., 1995, PASP, 107, 803
- Vaughan et al. (2016) Vaughan S., Uttley P., Markowitz A., Huppenkothen D., Middleton M., Alston W., Scargle J., Farr W., 2016, MNRAS, 461, 3145
- Villata & Raiteri (1999) Villata M., Raiteri C. M., 1999, A&A, 347, 30
- Wiita (2006) Wiita P. J., 2006, arXiv e-prints, pp astro–ph/0603728
- Wiita (2011) Wiita P. J., 2011, Journal of Astrophysics and Astronomy, 32, 147
- Xiong & Zhang (2014) Xiong D. R., Zhang X., 2014, Ap&SS, 352, 809
- Zanazzi & Lai (2019) Zanazzi J. J., Lai D., 2019, MNRAS, 487, 4965
- Zechmeister & Kürster (2009) Zechmeister M., Kürster M., 2009, A&A, 496, 577
- Zhou et al. (2018) Zhou J., Wang Z., Chen L., Wiita P. J., Vadakkumthani J., Morrell N., Zhang P., Zhang J., 2018, Nature Communications, 9, 4599
- del Palacio et al. (2019) del Palacio S., Bosch-Ramon V., Romero G. E., 2019, A&A, 623, A101
Appendix A Complete -ray LCs
This section includes the complete -ray LCs of the blazars studied in this paper.

### v3

@@ 4.2.2 External Jet Agents
Shock waves can also be generated by the interaction of external agents with the jet. One example is the penetration of dense clouds from the broad-line region into the jet. As these clouds interact with the relativistic flow, they induce strong shocks that efficiently accelerate particles, leading to enhanced synchrotron and inverse Compton emission, particularly in the -ray band (show_wave_clouds_palacio_2019). Another source of shocks can be the collision between the jet and stellar winds from massive stars within the host galaxy. These winds inject additional turbulence and inhomogeneity into the jet, generating shocks that modify the particle acceleration processes and impact the emission (shock_wave_stellar_wind_bicknell_2002).
4.2.3 Jet Precession
@@
Another phenomenon that can offer a compelling explanation for the dynamic evolution of the oscillation amplitude is the precession of the jet. In the scenario, changes in the Doppler boosting due to the variable orientation of the jet with respect to the line of sight cause changes in the amplitude of the oscillations over time (camenzind_jet). This is particularly evident when the precessing jet crosses critical angles that either amplify or attenuate the emitted radiation, creating a distinct trend in the LC (e.g., precession_chen_2024). The precession could be originated by multiple scenarios. One scenario could be the Lense-Thirring precession (e.g., franchini_lense; zanazzi_lense). This phenomenon is a relativistic effect caused by the dragging of spacetime due to the rotation of an SMBH. In blazars, the jet is typically aligned with the accretion disk’s angular momentum axis. However, if the accretion disk is misaligned with the SMBH spin axis, the frame-dragging effect from the rotating SMBH induces precession in the inner regions of the disk. This Lense-Thirring precession gradually changes the orientation of the jet, leading to observable variations in the direction of the jet over time. Alternatively, the presence of a binary SMBH system at the center of the blazar can be the origin of such precession (e.g., graham_binary; qian_precesion_binary). In the binary scenario, gravitational torques within the SMBH system induce precession, causing the relativistic jet of one SMBH to follow a slow, precessional path influenced by its companion. This gradual angular shift changes the jet angle relative to our line of sight, modifying the Doppler boosting effect (e.g., villata_helical_jet).
@@ 4.3 Interpretation
In this paper, we interpret the observed QPOs with multiplicative amplitudes as the result of a combination of different phenomena. First, we consider that QPOs arise from a plasmoid following a helical trajectory dictated by the magnetic field geometry. As discussed in 4, this motion alters the angle of our line of sight with respect to the jet, leading to periodic variations in the Doppler boosting factor. Additionally, the movement of the plasmoid can generate shock waves, which, in turn, induce fluctuations in the emitted flux. This effect can be perceived as an apparent change in the jet’s axis and, consequently, in our viewing angle (wiita_cells_2011), modulating the overall flux observed from the jet.
Combined with this, we also use the conclusions presented in rieger_2004 to explore how differential Doppler boosting, caused by the helical motion of emitting components within the jet, can lead to observable periodicity. This motion is nonballistic, meaning the radiating plasma elements follow curved trajectories rather than moving in straight lines. rieger_2004 identified three possible physical mechanisms for helical jet motion in blazars. The first is nonballistic motion driven by the orbital motion in an SMBH binary. The second involves either ballistic or nonballistic helical jet paths induced by jet precession, which may result from gravitational torques exerted by a secondary black hole or from accretion disk warping. The third mechanism attributes nonballistic helical motion to intrinsic jet instabilities. The first two scenarios are particularly viable for explaining 1 year.
@@ 4.3.1 Model expressions
as presented in otero_mwl. Additionally, the emitted flux can be expressed as
@@ 4.3.2 Flux Fit
Examining the shape of the oscillations (see Figure 1), we observe that their amplitude initially increases over time, reaching a peak, and subsequently decreases. Consequently, we substitute in Equation (4) with a linear function as =. This approach suggests that the amplitudes evolve over time through a multiplicative relationship (penil_2025). Therefore, our model accounts for the modulation introduced by the impact of the shock wave on the plasmoid’s emission. To extract the linear component of the flux, we employ a linear regression333We use the LinearRegression function of the Python package Scikit-learn, which is optimized specifically for linear fitting., characterizing the parameters of the linear trend444We use the seasonal_decompose function of the Python package Statsmodels..
@@
We then test the hypothesis of the LC being accurately described by this model through the R2 parameter, a statistical measure that assesses the goodness of fit of a regression model. The value of R2 ranges from 0 to 1 (or 0 to 100 in percentage), where the higher values denote a better fit and suggest that the model can better explain the variations in the data. In hair_r2_2011, three different levels of R2—25%, 50%, and 75%—are outlined and categorized as weak, moderate, and substantial, respectively to evaluate the accuracy of a fit.
@@ 5 Results and Discussion
The Lorentz factor, , generally falls within the range of approximately 10 to 20 for FSRQs (gamma_fsrq_ghisellini), often reaching values (xiong2014). In this study, we adopt (ghisellini_gamma; sarkar_curved_jet). The parameter is directly related to through the expression , which yields . The pitch angle, , is another variable with a potentially broad range, typically between and (e.g., jorstad_pitch_angle; butuzova_pitch_angle). However, owing to the lack of observational constraints in the literature, we adopt a pitch angle value of as done in previous studies (sarkar_curved_jet; anuvab_curve_jet).
For the values of the period , the linear slope, the intrinsic spectral index , and the Doppler boosting factor , we conduct an exploratory process to achieve the optimal flux fit modeled by Equation (4) for each LC by maximizing the R2 value. For the periods and slopes, this optimization spans the range of periods and slopes determined through the methodology outlined in . The resulting values are summarized in Table 2.
Regarding , spectral_index_fsrq_gamma conducted a comprehensive analysis of blazars detected by Fermi-LAT, reporting a mean photon spectral index of 1.4. Consistently, we explore values for in the range of 1.2, 1.3, 1.4, 1.5, and 1.6 to better understand the variations in the spectral behavior.
Finally, for the Doppler boosting factor , we explore different values, from to . When the emission in the jet is primarily due to synchrotron radiation, . This lower value reflects the isotropic nature of synchrotron emission in the rest frame of the jet, typically prominent in radio to optical wavelengths and causing moderate flux variability in these bands (urry_n). For rays, which are often produced via inverse Compton scattering, tends to be higher, typically 3 or 4 (dermer_doopler). This increase occurs because inverse Compton scattering — and especially external Compton scattering, where seed photons from outside the jet are scattered — is more sensitive to the Doppler factor. This phenomenon is particularly dominant in FSRQs, which constitute the objects in our sample. Finally, is particularly relevant in scenarios where short-lived or transient events are observed, suggesting that a localized, highly energetic region within the jet contributes to extreme, rapid flux variations characteristic of the FSRQ (dermer_doopler). The specific values of the periods, slopes, , and that maximize the R2 value are displayed in Figure 2 and Table 2, highlighting the most accurate fits achieved for each case.
@@
As discussed in 4.3.2, we perform a detailed analysis of the QPOs by separately examining the increasing and decreasing segments of the LC, characterizing potential variations in amplitude or period between the rising and falling phases. We apply Equation (4) to model each segment, optimizing the variable values to maximize R2.
@@
To illustrate the evolution of , we present Figure 3, which depicts its temporal variation based on different parametrization results for the three blazars. This visualization shows how evolves over time, helping to assess the theoretical scenarios considered for interpreting the observed flux variations. In the specific case of 4C +31.03, we observe that the viewing angle evolution shows a discontinuity of 30% between the increasing and decreasing segments of the LC. Both the low R2 value and this discontinuity suggest potential inconsistencies in modeling the decreasing segment. The heightened variability within this portion of the LC could be impacting the analysis accuracy, as fluctuations in the -ray emission introduce complexities that may not align well with the current model assumptions. These results raise questions about the suitability of the selected theoretical model in capturing the transient QPO behavior of 4C +31.03, suggesting that an alternative approach may be necessary to better understand the emission mechanisms in this blazar. One possible scenario is that recurring boosts in the turbulent flow behind a propagating shock generate the QPO. Due to the stochastic nature of turbulence, variations in its properties can cause fluctuations in both the QPO period and the emitted flux (wiita_cells_2011), leading to changes in the observed period, as seen in the case of 4C +31.03 (see Table 2).
@@ 5.2 MG1 J123931+0443
The flux fits for both segments of the LC separated by the flare located at 56400 MJD present consistent results, with compatible periods and trends observed across both segments, as shown in Table 2 and Figure 2. Additionally, Figure 3 reveals no significant discontinuity, with only a minor variation of 6.6%, which can be attributed to uncertainties introduced by the flaring state during parameter estimation. In fact, from the QPO analysis, we observe that this error in the is consistent within errors at a 1 confidence level. These findings reinforce the accuracy of the chosen emission model, supporting it as a plausible scenario to explain the QPO observed in MG1 J123931+0443.
As discussed in 4, several theoretical scenarios can provide explanations for the observed QPO. rieger_2004 suggests that periodicity driven by differential Doppler boosting could result from (1) orbital motion in a close binary SMBH, (2) Newtonian precession within a close SMBH, or (3) internal jet rotation. Among these, Newtonian-driven precession appears particularly viable for explaining periods of year, which aligns with the observed period in MG1 J123931+0443, 285 days. If the period observed in this object is associated with the presence of a binary SMBH, we can relate with the intrinsic orbital period (), as
where is the redshift of the source, and the motion of the plasma is considered as nonballistic and helical, driven by the binary system’s orbital dynamics (rieger_2004). A nonballistic motion occurs when the trajectory of the plasma is continuously influenced by external forces, such as the jet’s environment or the central engine of the blazar (rieger_2004). For an observed period days and (see Table 1), the calculated orbital period is years assuming a .
Using this orbital period, we estimate the mass of the primary SMBH, , based on the formulation provided by ostorero_2004,
where the mass ratio typically ranges from to (ostorero_2004). Substituting the orbital period, we estimate the primary SMBH’s mass to range from 7.7 to 1.2, consistent with reasonable mass estimates for SMBHs (liang_2003). mg1_J123931_0443_massbh estimated the SMBH mass of this blazar using various techniques, obtaining a range of 2.1-9.7, consistent with the range of values reported here.
@@
is the gravitational constant and is the mass of the secondary SMBHs. This yields an orbital separation in the range of 8.9 to 1.8 parsecs. These values are consistent with those expected for binary SMBH systems, typically estimated to lie between and pc (bogdanovic_2008; colpi_2014).
If the primary SMBH emits a relativistic jet, its behavior may be influenced by the binary system’s orbital dynamics. The orbital radius, defined as the distance between the primary SMBH and the system’s barycenter (ostorero_2004), is calculated as:
@@ 5.3 PKS 1622253
This blazar exhibits a high level of consistency between the estimated period and trend for both segments of the LC, further supported by the minimal discontinuity in , measured at 3.6%. The observed period, days, also suggests that Newtonian precession due to the interaction within a binary SMBH system could be a plausible physical scenario, as in the case of MG1 J123931+0443.
We characterize the candidate binary SMBH system of PKS 1622253. Using the redshift value of and the aforementioned expressions. We derive an orbital period of 117 years. The estimation of the primary SMBH’s mass ranges from 2.0 to 2.2, consistent with the mass of 4.0 reported by gamma_fsrq_ghisellini. The orbital separation ranges of 1.8 to 3.7 parsecs. The orbital radius ranges between 9.2 and 3.7 parsecs.
@@ 5.4 Precession Model
An alternative scenario could be the precessing jet model, which was mentioned in 4. In this scenario, the relativistic jet emitted by a blazar undergoes precession, causing the jet’s orientation to change over time. This variation alters the angle between the jet and our line of sight, influencing the observed emission due to changes in the relativistic beaming effects, potentially leading to the observed periodicity in the blazar’s LCs (camenzind_jet).
@@ 6 Summary
The QPO features detected in PKS 1622253 and MG1 J123931+0443 show a combination of year-like periodicities and smooth linear amplitude modulation, which align with the expectations for differential Doppler boosting induced by the presence of a binary SMBH system. In contrast, 4C +31.03 exhibits less coherence, more typical of stochastic variability. Therefore, PKS 1622253 and MG1 J123931+0443 emerge as binary SMBH candidates. However, given the relatively low significance derived from our transient QPO analysis, these results should be interpreted cautiously as tentative hints rather than definitive evidence. Additional studies and observations will be essential to further test and validate this interpretation. Despite these limitations, the consistency between the theoretical model and the observed -ray LCs highlights the plausibility that binary SMBH systems could be involved in producing QPOs in blazars, offering potentially valuable insights into the nature of periodic variability in these sources.
Acknowledgements
@@ Acknowledgements.
P.P. and M.A. acknowledge funding under NASA contract 80NSSC20K1562. J.O.-S. acknowledges founding from the Istituto Nazionale di Fisica Nucleare (INFN) Cap. U.1.01.01.01.009. This work was supported by the European Research Council, ERC Starting grant MessMapp, S.B. Principal Investigator, under contract no. 949555, and by the German Science Foundation DFG, research grant “Relativistic Jets in Active Galaxies” (FOR 5195, grant No. 443220636).Appendix A Complete -ray LCs
This section includes the complete -ray LCs of the blazars studied in this paper, shown in Figure A1.